OpenStudy (anonymous):

If tan θ =(√3)/3, 0◦< θ < 360◦, find the smaller θ in degrees. Answer in units of degrees. and radians

OpenStudy (anonymous):



OpenStudy (turingtest):

tan θ =(√3)/3\[\theta = \tan^{-1}(\sqrt3/3)=0.523rad\]\[\pi rad=180\deg\] 0.532rad(180deg/pi rad)=30 degrees

OpenStudy (anonymous):

find the larger θ in degrees. Answer in units of degrees. and radians

OpenStudy (anonymous):

help me please. i wanna see how its solved.